A cellular phone company has the following production function for a smart phone: p(x, y) =

50x

2

3 y

1

3 where p is the number of units produced with x units of labor and y units of capital. a)

Find the number of units produced with 125 units of labor and 64 units of capital. b) Find the

marginal productivities (Hints: Partial derivatives). c) Evaluate the marginal productivities at

x = 125 and y = 64.



1
Expert's answer
2021-10-27T11:28:15-0400

Solution:

Let the production function be "p(x,y)=50x^{2\/3}y^{1\/3}"

(a)

"p(125,64)=50(125)^{2/3}(64)^{1/3}
\\ =50(25)(4)
\\=5000\ units"

(b)

Marginal productivity of labor"=\dfrac{\partial p}{\partial x}=p_x"

"=50\cdot\dfrac23x^{-1/3}y^{1/3}=\dfrac{100y^{1/3}}{3x^{1/3}}"

Marginal productivity of capital"=\dfrac{\partial p}{\partial y}=p_y"

"=50\cdot\dfrac13x^{2/3}y^{-2/3}=\dfrac{50x^{2/3}}{3y^{2/3}}"

(c)

Marginal productivity of labor"=p_x(125,64)"

"=\dfrac{100(64)^{1/3}}{3(125)^{1/3}}
\\=\dfrac{100\times 4}{3\times 5}
\\=26 \dfrac23"

Marginal productivity of capital"=p_y(125,64)"

"=\dfrac{50(125)^{2/3}}{3(64)^{2/3}}
\\=\dfrac{50\times25}{3\times16}
\\=26 \dfrac1{24}"